4. Consider a bag that contains 221 coins of which 4 are rare Indian pennies. For the given pair of events A and​ B, complete parts​ (a) and​ (b) below.  

​A: When one of the 221 coins is randomly​ selected, it is one of the 4 Indian pennies

​B: When another one of the 221 coins is randomly selected​ (with replacement), it is also one of the 4 Indian pennies. a. Determine whether events A and B are independent or dependent.

b. Find​ P(A and​ B), the probability that events A and B both occur.

b.The probability that events A and B both occur is____.

​(Round to six decimal places as​ needed.)

5. For the given pair of events A and​ B, complete parts​ (a) and​ (b) below.

​A: A marble is randomly selected from a bag containing 10 marbles consisting of 1​ red, 6 ​blue, and 3 green marbles. The selected marble is one of the green marbles.

​B: A second marble is selected and it is the 1 red marble in the bag.

a. Determine whether events A and B are independent or dependent.​ (If two events are technically dependent but can be treated as if they are independent according to the​ 5% guideline, consider them to be​ independent.)

b. Find​ P(A and​ B), the probability that events A and B both occur.

b. The probability that events A and B both occur is____.

​(Round to four decimal places as​ needed.)
Answer #1

4)

a)

as experiment is done with replacement

The two events are independent because the occurrence of one does not affect the probability of the occurrence of the other

b)

P(A)=4/221

P(B with replacement) = 4/221

P(A and B)=4/221*4/221=0.000328

----------------------------------

please note that only one ques per post can be answered as per HOMEWORKLIB RULES